for what value(s) of 'n' does the function f(x)=16x^2-8x+n have exactly one zero?
Answer by ikleyn(52778)

.

The discriminant of the quadratic function must be equal to zero


    d = b^2 - 4ac = (-8)^2 - 4*16*n = 0,

or

    64 = 64*n

     n = 64/64 = 1.     ANSWER

Solved.